OpenStudy (anonymous):

Express the complex number in trigonometric form. 3 - 3i

zepdrix (zepdrix):

To find your angle, \[\Large\rm \theta=\arctan \frac{y}{x}\]Plug in the x and y,\[\Large\rm \theta=\arctan \frac{-3}{3},\qquad\qquad\to\qquad \qquad \tan \theta=-1\] Then to find the size of your radius,\[\Large\rm r=|3-3\mathcal i|=\sqrt{3^2+3^2}\]Then you can write it in this form:\[\Large\rm r(\cos \theta+ i \sin \theta)\]

OpenStudy (anonymous):

none of that matches any of my answer choices

zepdrix (zepdrix):

I didn't give you an answer, I gave you the steps to get your answer.

OpenStudy (anonymous):

well it does not make sense so its not helping